# Distributional Random Forests for Complex Survey Designs on Reproducing Kernel Hilbert Spaces

### v1

@@ 4.2 Computation Details
We run the SDRF and DRF ([21]) algorithms with the above data generating process for each of the finite-population size $N\in\{4000,5000,15000,22500\}$ and the number of trees $B\in\{10,30,70,200\}$. For each of the scenario, 200 seeds were run. We let $p=3$ to align with the case study (Section 5). As $N$ increases, we also increase the total number of strata $H$ and the number of PSUs per stratum $N_{\mathrm{psu},h}$. The expected sample size per stratum (and hence the overall sample size) increases with $N$. The sampling mechanism (PPS at stage 1 and SRSWOR at stage 2 with fixed rates) is kept fixed in the values of $N$, which is consistent with the asymptotics of the increase in the population survey.
@@
$\Psi(\mathbf{x}=(x_{1},\dots,x_{p}))\;=\;\mathbb{E}\!\left[\,Y_{1}\mid X=\mathbf{x}_{1}\,\right],$
@@
For both random-forest estimators, we use a Gaussian kernel on the outcome space with bandwidth set by the median of pairwise Euclidean distances between the observed $Y$’s. The maximum tree depth is fixed at $8$. The minimum node size is set to $\max\!\left\{\,20,\;\left\lceil\sqrt{n_{s}}\,\right\rceil\right\}$, where $n$ denotes the size of the training-sample. We cap effective sampling weights by $\lambda_{\max}\;=\;5.5\times\frac{\max_{i}w_{i}}{\min_{i}w_{i}}$, where $w_{i}$ are the survey weights. Although additional hyperparameter tuning could further improve performance, we keep the settings minimal because the untuned configurations already reveal the main trends across $N$ and $B$.
We do not compare against alternative survey-calibrated distributional methods beyond SDRF because, to our knowledge, such methods are not currently available in the survey-sampling literature. (The unweighted DRF serves as the i.i.d. baseline and is a logical competitor). All code is implemented under R, version 4.4.2 [73].
@@ 4.3 Results
Table 1 reports the mean (SD) of the MMD between the true finite population conditional distribution $P(Y\mid X)$ and its estimate $\widehat{P}(Y\mid X)$ from SDRF. The mean MMD decreases as the finite population size $N$ increases, while increasing the number of trees $B$ primarily reduces the SD (stabilizing forest performance) with comparatively smaller improvements in the mean and diminishing returns beyond moderate $B$. Overall, these finite-sample patterns are consistent with the convergence behavior asserted in Theorem 3.4.
Table 1: Mean (SD) of $100\times d_{\mathbf{k}}(\widehat{P}(Y\mid X),\;P^{N}(Y\mid X))$, where $\widehat{P}(Y\mid X)$ is the distribution estimate from SDRF, across finite population size $N$ and number of trees $B$.
NN $B=10$ $B=30$ $B=70$ $B=200$
4,000 19.60 (2.52) 18.60 (2.27) 18.30 (2.19) 18.30 (2.10)
5,000 19.30 (2.49) 18.40 (2.14) 18.00 (2.14) 18.10 (2.07)
15,000 17.30 (1.86) 16.80 (1.44) 16.70 (1.30) 16.80 (1.23)
22,500 17.40 (1.52) 16.80 (1.23) 16.60 (1.16) 16.70 (1.08)
Table 2: Root mean squared error (RMSE) of the conditional mean estimator $\widehat{\mathbb{E}}[Y_{1}|X_{1}]$ averaged over $X_{1}$ for Survey-DRF (SDRF) and benchmark DRF algorithms as $N$ and the number of trees grow.
@@ 5 Biomedical Public Health Case Study
This section demonstrates the practical relevance of survey-calibrated distributional random forests (SDRF) for performing data analysis tasks in nationally representative studies. We use data from the National Health and Nutrition Examination Survey (NHANES), which employs a multistage, stratified, cluster sampling design to represent the U.S. civilian, noninstitutionalized population: primary sampling units (PSUs) are selected at the county level, followed by smaller geographic areas and then individuals. All analyzes performed incorporate survey weights and design features provided by the Centers for Disease Control and Prevention (CDC) [^1] https://wwwn.cdc.gov/nchs/nhanes/Default.aspx . More specifically, in this paper, we focus on the 2011–2012 NHANES cycle and apply an SDRF algorithm that sampling mechanism of the NHANES study during bagging/subsampling and throughout tree construction (e.g. split selection and node estimation).
The objective of the statistical analysis is to characterize the conditional joint distribution of two glucose biomarkers used to diagnose and monitor the progression of diabetes, fasting plasma glucose (FPG) and glycated hemoglobin (HbA1c), given demographic and anthropometric covariates: sex, age, and body mass index (BMI). Our primary clinical goal is to quantify how the population distribution of this bivariate outcome varies between these characteristics, to support the detection of diabetes, the evaluation of glycemic-control, and healthcare planning through the trends observed in the subgroups of the population. A secondary clinical objective is to evaluate whether the observed heterogeneity motivates individual diabetes diagnostic thresholds for FPG and HbA1c, in contrast to fixed cutoffs proposed by medical associations such as the American Diabetes Association (ADA) [2], which ignore specific individual characteristics. For analytical purposes, a total of $N=2732$ individuals with and without diabetes were included in the analysis.
Let us go to define mathematically the problem. Let the clinical outcome be $Y=(Y_{1},Y_{2})^{\top}$, where $Y_{1}$ denotes fasting plasma glucose (FPG) and $Y_{2}$ denotes HbA1c. Let the covariate vector be $X=(X_{1},X_{2},X_{3})^{\top}$, where $X_{1}=\mathrm{Age}$ and $X_{2}=\mathrm{BMI}$ are continuous, and $X_{3}=\mathrm{Sex}\in\{0,1\}$ is binary with $0=\text{male}$ and $1=\text{female}$.
From an empirical perspective, we run SDRF and estimate the conditional distribution of the bivariate result $Y=(\mathrm{FPG},\mathrm{HbA1c})^{\top}$ given covariates $X=\mathbf{x}$. Specifically, we computed the estimated conditional CDF
@@
where inequality $Y\leq y$ is understood componentwise. For representation purposes, we evaluate $\hat{F}(\cdot\mid\mathbf{x})$ on a grid of covariate combinations:
@@
Let $\hat{\omega}_{i}(x)$ denote the SDRF distribution-regression weights for observation $i\in[N]$ at covariate value $x\in\mathcal{X}$. Using these weights, we compute the estimated conditional mean and covariance,
$\hat{\mu}(x)=\sum_{i=1}^{N}\hat{\omega}_{i}(x)\,Y_{i}\quad\text{and}\quad\hat{\Sigma}(x)=\sum_{i=1}^{N}\hat{\omega}_{i}(x)\,\bigl(Y_{i}-\hat{\mu}(x)\bigr)\bigl(Y_{i}-\hat{\mu}(x)\bigr)^{\top}.$
We then define a conditional Mahalanobis score (weighted) for any pair $(x,y)\in\mathcal{X}\times\mathcal{Y}$ by
$S(x,y)=\bigl(y-\hat{\mu}(x)\bigr)^{\top}\hat{\Sigma}(x)^{-1}\bigl(y-\hat{\mu}(x)\bigr).$
Let $\{(x_{a},y_{a})\}_{a=1}^{N}$ denote the observed random sample and let $\hat{q}_{1-\alpha}$ be the empirical $(1-\alpha)$ quantile of the sample $\{S(x_{a},y_{a})\}_{a=1}^{N}$ (using the survey weights).
In the spirit of [53] (for finite-dimensional Euclidean responses) and [57] (for metric space outcomes), we define the tolerance region $(1-\alpha)$ at the covariate value $x\in\mathcal{X}$ as
$\hat{C}^{\alpha}(x)=\bigl\{\,y\in\mathcal{Y}:\;S(x,y)\leq\hat{q}_{1-\alpha}\,\bigr\}.$
Compared to the natural competitor [53], our approach is more flexible: it requires a single model, avoids user-chosen smoothing parameters due to the use of multiple local polynomial regression, and enables valid inference under the complex design of the NHANES survey.
@@
We discuss the statistically supported variable by variable. The estimated joint distributions for women and men are similar in shape; however, women exhibit more extreme observation and a higher prevalence of elevated fasting plasma glucose (FPG), especially at higher $\alpha\in(0,1)$ levels. With increasing age, both FPG and HbA1c shift upward, indicating poorer glycemic control at older ages; this effect is most pronounced and consistent near the $0.9$ quantile. A higher body mass index (BMI) is moderately associated with higher glucose levels: people with $\mathrm{BMI}>30\penalty 10000\ \mathrm{kg}/\mathrm{m}^{2}$ show higher values than those with $\mathrm{BMI}\in[20,30]\penalty 10000\ \mathrm{kg}/\mathrm{m}^{2}$. Between $30$ and $40\penalty 10000\ \mathrm{kg}/\mathrm{m}^{2}$, the tolerance regions change little between the quantile levels. In general, glucose control differs between overweight/obesity individuals and those with a BMI in the normal range.
From these results, two implications follow: (i) glycemic control in the U.S. population varies significantly between demographic and anthropometric individual characteristics, with consequences for public-health planning and political interventions; and (ii) diagnostic cutoff points for diabetes based on FPG and HbA1c may warrant adjustment for at least age and possibly BMI, rather than relying on the same rules for all population in line with [8] . From a modeling and analysis perspective, SDRF provides a principled framework to define reference values for specific subpopulations, reclassify individuals based on biomarker profiles, and forecast public–health risks leveraging the full conditional distribution. In contrast, approaches that model only the conditional mean can miss clinically important features—such as quantiles, dispersion, and tail behavior, leading to less informative inform to perform the decision making.
@@ 6 Discussion
We propose a random forest tailored to complex survey designs, in which the bagging step explicitly mimics the sampling design (e.g., strata, primary sampling units, and weights), inspired by recent progress on bootstrap methods for complex designs with rate improvements [95]. To the best of our knowledge, random forests and other tree ensembles have not previously been adapted to complex survey data, filling an important methodological gap. From an estimation and applied data analysis perspective, we introduce what is, to our knowledge, the first general framework for complex designs to estimate conditional distributions for multivariate outcomes. Our approach leverages reproducing-kernel Hilbert spaces (RKHS) and distances between kernel mean embeddings to estimate the full conditional distribution. This representation enables straightforward derivation of population quantities (e.g., quantiles, ratios) and supports downstream tasks–including two-sample testing, selection of distributional variables, and semi-supervised clustering—via a forest-based kernel mean embedding approach [7, 67, 79, 37].
SDRF framework provides a powerful and flexible tool for biomedical studies conducted under complex survey designs, as illustrated by the NHANES case study reported here, studies that are essential for optimizing public-health decisions. Our machine learning algorithm for estimating the conditional distribution of multivariate outcomes represents a substantial modeling advance, enabling investigators to address clinically relevant questions in these cohorts with direct implications for health policy, as demonstrated in our case study illustration.
@@
In future work, we will continue to investigate the survey-design aspects of our method. In particular, we will study the upward bias that can arise in ratio-type estimators when the denominator (here, the leaf size or subpopulation size) must be estimated because the true leaf sizes are not available from the finite population or census. Extending SDRF for outcome-dependent sampling is another direction.

### v2

@@ 4.2 Computation Details
We run the SDRF and DRF (cevid_michel_meinshausen_bühlmann_2022) algorithms with the above data generating process for each of the finite-population size $N\in\{4000,5000,15000,22500\}$ and the number of trees $B\in\{10,30,70,200\}$. For each of the scenario, 200 seeds were run. We let $p=3$ to align with the case study (Section 5). As $N$ increases, we also increase the total number of strata $H$ and the number of PSUs per stratum $N_{\mathrm{psu},h}$. The expected sample size per stratum (and hence the overall sample size) increases with $N$. The sampling mechanism (PPS at stage 1 and SRSWOR at stage 2 with fixed rates) is kept fixed in the values of $N$, which is consistent with the asymptotics of the increase in the population survey.
@@
$\Psi(\mathbf{x}=(\mathbf{x}_{1},\dots,\mathbf{x}_{p}))\;=\;\mathbb{E}\!\left[\,Y_{1}\mid X=\mathbf{x}_{1}\,\right],$
@@
For both random-forest estimators, we use a Gaussian kernel on the outcome space with bandwidth set by the median of pairwise Euclidean distances between the observed $Y$’s. The maximum tree depth is fixed at $8$. The minimum node size is set to $\max\!\left\{\,20,\;\left\lceil\sqrt{n_{s}}\,\right\rceil\right\}$, where $n$ denotes the size of the training-sample. We cap node–level effective sampling weights by $\lambda_{\max}\;=\;5.5\times\frac{\max_{i}w_{i}}{\min_{i}w_{i}}$, where $w_{i}$ are the survey weights. Although additional hyperparameter tuning could further improve performance, we keep the settings minimal because the untuned configurations already reveal the main trends across $N$ and $B$.
We do not compare against alternative survey-calibrated distributional methods beyond SDRF because, to our knowledge, such methods are not currently available in the survey-sampling literature. (The unweighted DRF serves as the i.i.d. baseline and is a logical competitor). All code is implemented under R, version 4.4.2 R_cite.
@@ 4.3 Results
Table 1 reports the mean (SD) of the MMD between the true finite population conditional distribution $P_{S^{N}|\omega}(Y\mid X)$ and its estimate $\widehat{P}_{S^{N}|\omega}(Y\mid X)$ from SDRF. The mean MMD decreases as the finite population size $N$ increases, while increasing the number of trees $B$ primarily reduces the SD (stabilizing forest performance) with comparatively smaller improvements in the mean and diminishing returns beyond moderate $B$. Overall, these finite-sample patterns are consistent with the convergence behavior asserted in Theorem 3.4.
Table 1: Mean (SD) of $100\times d_{\mathbf{k}}(\widehat{P}_{S^{N}|\omega}(Y\mid X),\;P_{S^{N}|\omega}(Y\mid X))$ across finite population size NN (hence different average sample sizes $\bar{n}_{s}$) and the number of trees $B$.
NN $\bar{n}_{s}$ $B=10$ $B=30$ $B=70$ $B=200$
4,000 320 19.60 (2.52) 18.60 (2.27) 18.30 (2.19) 18.30 (2.10)
5,000 400 19.30 (2.49) 18.40 (2.14) 18.00 (2.14) 18.10 (2.07)
15,000 1,200 17.30 (1.86) 16.80 (1.44) 16.70 (1.30) 16.80 (1.23)
22,500 1,800 17.40 (1.52) 16.80 (1.23) 16.60 (1.16) 16.70 (1.08)
Table 2: Root mean squared error (RMSE) of the conditional mean estimator $\widehat{\mathbb{E}}[Y_{1}|X_{1}]$ averaged over $X_{1}$ for SDRF and benchmark DRF algorithms as finite population size $N$ and the number of trees $B$ grow.
@@ 5 Biomedical Public Health Case Study
This section demonstrates the practical relevance of survey-calibrated distributional random forests (SDRF) for performing data analysis tasks in nationally representative studies. We use data from the National Health and Nutrition Examination Survey (NHANES), which employs a multistage, stratified, cluster sampling design to represent the U.S. civilian, noninstitutionalized population: primary sampling units (PSUs) are selected at the county level, followed by smaller geographic areas and then individuals. All analyses performed incorporate survey weights and design features provided by the Centers for Disease Control and Prevention (CDC) [^1] https://wwwn.cdc.gov/nchs/nhanes/Default.aspx . More specifically, in this paper, we focus on the 2011–2012 NHANES cycle and apply an SDRF algorithm that incorporates the sampling mechanism of the NHANES study during bagging/subsampling and throughout tree construction (e.g. split selection and node estimation).
The objective of the statistical analysis is to characterize the conditional joint distribution of two glucose biomarkers used to diagnose and monitor the progression of diabetes, fasting plasma glucose (FPG) and glycated hemoglobin (HbA1c), given demographic and anthropometric covariates: sex, age, and body mass index (BMI). Our primary clinical goal is to quantify how the population distribution of this bivariate outcome varies between these characteristics, to support the detection of diabetes, the evaluation of glycemic-control, and healthcare planning through the trends observed in the subgroups of the population. A secondary clinical objective is to evaluate whether the observed heterogeneity motivates individual diabetes diagnostic thresholds for FPG and HbA1c, in contrast to fixed cutoffs proposed by medical associations such as the American Diabetes Association (ADA) american20252, which ignore specific individual characteristics. For analytical purposes, a total of $N=2732$ individuals with and without diabetes were included in the analysis.
We now articulate the analysis goal mathematically. Let the clinical outcome be $Y=(Y_{1},Y_{2})^{\top}$, where $Y_{1}$ denotes fasting plasma glucose (FPG) and $Y_{2}$ denotes HbA1c. Let the covariate vector be $X=(X_{1},X_{2},X_{3})^{\top}$, where $X_{1}=\mathrm{Age}$ and $X_{2}=\mathrm{BMI}$ are continuous, and $X_{3}=\mathrm{Sex}\in\{0,1\}$ is binary with $0=\text{male}$ and $1=\text{female}$.
From an empirical perspective, we run SDRF and estimate the conditional distribution of the bivariate result $Y=(\mathrm{FPG},\mathrm{HbA1c})^{\top}$ given covariates $X=\mathbf{x}$. Specifically, we compute the estimated conditional CDF
@@
where inequality $Y\leq y$ is understood component–wise. For representation purposes, we evaluate $\hat{F}(\cdot\mid\mathbf{x})$ on a grid of covariate combinations:
@@
Let $\hat{\omega}_{i}(\mathbf{x})$ denote the SDRF distribution-regression weights for observation $i\in[N]$ at covariate value $\mathbf{x}\in\mathcal{X}$. Using these weights, we compute the estimated conditional mean and covariance,
$\hat{\mu}(\mathbf{x})=\sum_{i=1}^{N}\hat{\omega}_{i}(\mathbf{x})\,Y_{i}\quad\text{and}\quad\hat{\Sigma}(\mathbf{x})=\sum_{i=1}^{N}\hat{\omega}_{i}(\mathbf{x})\,\bigl(Y_{i}-\hat{\mu}(\mathbf{x})\bigr)\bigl(Y_{i}-\hat{\mu}(\mathbf{x})\bigr)^{\top}.$
We then define a conditional Mahalanobis score (weighted) for any pair $(\mathbf{x},y)\in\mathcal{X}\times\mathcal{Y}$ by
$S(\mathbf{x},y)=\bigl(y-\hat{\mu}(\mathbf{x})\bigr)^{\top}\hat{\Sigma}(\mathbf{x})^{-1}\bigl(y-\hat{\mu}(\mathbf{x})\bigr).$
Let $\{(\mathbf{x}_{a},y_{a})\}_{a=1}^{N}$ denote the observed random sample and let $\hat{q}_{1-\alpha}$ be the empirical $(1-\alpha)$ quantile of the sample $\{S(\mathbf{x}_{a},y_{a})\}_{a=1}^{N}$ (using the survey weights).
In the spirit of lado2021 (for finite-dimensional Euclidean responses) and lugosi2025conformal (for metric space outcomes), we define the tolerance region $(1-\alpha)$ at the covariate value $\mathbf{x}\in\mathcal{X}$ as
$\hat{C}^{\alpha}(\mathbf{x})=\bigl\{\,y\in\mathcal{Y}:\;S(\mathbf{x},y)\leq\hat{q}_{1-\alpha}\,\bigr\}.$
A natural competitor to SDRF in this setting is the method of Lado et al. (2021) lado2021. That approach fits multiple univariate local polynomial regressions to estimate the conditional mean, and then uses residual-based estimates to approximate the conditional covariance and construct a $(1-\alpha)$ tolerance region. Because it proceeds marginally via separate univariate fits, it is unlikely to be fully efficient for joint estimation. Moreover, the required selection of smoothing parameters is performed without an explicit global criterion, which can lead to suboptimal choices. Finally, the method does not incorporate survey weights during estimation. In contrast, the SDRF approach is more flexible: it requires only a single model, avoids the user-chosen smoothing parameters required by multiple local polynomial regressions, and enables valid inference under the complex design of the NHANES survey.
@@
We discuss the statistical results variable by variable. The estimated joint distributions for women and men are similar in shape; however, women exhibit more extreme observations and a higher prevalence of elevated fasting plasma glucose (FPG), especially at higher $\alpha\in(0,1)$ levels. With increasing age, both FPG and HbA1c shift upward, indicating poorer glycemic control at older ages; this effect is most pronounced and consistent near the $0.9$ quantile. A higher body mass index (BMI) is moderately associated with higher glucose levels: people with $\mathrm{BMI}>30\penalty 10000\ \mathrm{kg}/\mathrm{m}^{2}$ show higher values than those with $\mathrm{BMI}\in[20,30]\penalty 10000\ \mathrm{kg}/\mathrm{m}^{2}$. Between $30$ and $40\penalty 10000\ \mathrm{kg}/\mathrm{m}^{2}$, the tolerance regions change little between the quantile levels. In general, glucose control differs between overweight/obesity individuals and those with a BMI in the normal range.
From these results, two implications follow: (i) glycemic control in the U.S. population varies significantly between demographic and anthropometric individual characteristics, with consequences for public-health planning and potential policy interventions; and (ii) diagnostic cutoff points for diabetes based on FPG and HbA1c may warrant adjustment for at least age and possibly BMI, rather than relying on the same rules for all population in line with Berkovic2025 . From a modeling and analysis perspective, SDRF provides a principled framework to define reference values for specific subpopulations, reclassify individuals based on biomarker profiles, and forecast public–health risks leveraging the full conditional distribution. In contrast, approaches that model only the conditional mean can miss clinically important features—such as quantiles, dispersion, and tail behavior, leading to lower richness of information for decision making.
@@ 6 Discussion
We propose a random forest tailored to complex survey designs, in which the bagging step explicitly mimics the sampling design (e.g., strata, primary sampling units, and weights), inspired by recent progress on bootstrap methods for complex designs with rate improvements wang2022bootstrap. To the best of our knowledge, random forests and other tree ensembles have not previously been adapted to complex survey data, thus our proposed approach fills an important methodological gap. From an estimation and applied data analysis perspective, we introduce what is, to our knowledge, the first general framework for complex designs for estimating conditional distributions for multivariate outcomes. Our approach leverages reproducing-kernel Hilbert spaces (RKHS) and distances between kernel mean embeddings to estimate the full conditional distribution. This representation enables straightforward derivation of population quantities (e.g., quantiles, ratios) and supports downstream tasks–including two-sample testing, selection of distributional variables, and semi-supervised clustering—via a forest-based kernel mean embedding approach benard2024mmd; matabuena2025kernel; scornet2016random; gretton2012kernel.
The SDRF framework provides a powerful and flexible tool for research studies conducted under complex survey designs, as illustrated by the NHANES case study reported here. Such studies are essential for optimizing public-health decisions. Our machine learning algorithm for estimating the conditional distribution of multivariate outcomes represents a substantial modeling advance, enabling investigators to address clinically relevant questions in these cohorts with direct implications for health policy, as demonstrated in our case study illustration.
@@
In future work, we will continue to investigate the survey-design aspects of our method. In particular, we will study the upward bias that can arise in ratio-type estimators when the denominator (here, the leaf size or subpopulation size) must be estimated because the true leaf sizes are not available from the finite population or census. Extending SDRF for outcome-dependent sampling is another potentially interesting direction.